OpenStudy (anonymous):

Use the multiplication principle of equality to eliminate the fractions, then solve. 6/5s+4/5=5/2+3/5s

OpenStudy (anonymous):

6/5s-3/5s=5/2-4/5

OpenStudy (anonymous):

(6-3)/5s=(25-8)/10

OpenStudy (anonymous):

3/5s=17/10

OpenStudy (anonymous):

30=85s

OpenStudy (anonymous):

s=30/85

OpenStudy (anonymous):

s=6/17

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome